Suppose that scores on a national entrance exam are normally distributed with mean 1000 and standard deviation 100.Which of the following is false?
1
Expert's answer
2020-09-14T15:23:23-0400

Given :

mean =1000

sd=100

Solution :


Variance = 1000 is false because variance = "(SD)^2=(100)^2=10000"